OpenStudy (anonymous):

derivative of f(x) = (x-2)^(1/3) X (x+1)^(2/3) @Mathematics

OpenStudy (anonymous):

yikes

OpenStudy (cwrw238):

use product rule

OpenStudy (cwrw238):

f(x) = (x-2)^(1/3 , f'(x) = (1/3) (x-2)^ (-2/3) g(x) = (x+1)^(2/3) , g'(x) = (2/3) (x+1)^(-1/3) and reqd. derivative = f(x)g'(x) + f'(x)g(x)

OpenStudy (shadowfiend):

Please don't tag OpenStudy Feedback in questions that aren't feedback on the site. Thanks!